Update #1:  The Metop-B HIRS was placed into High Power mode on May 28, 2020 (20/149) at 1526 UTC.  The Filter Motor current rose to 446mA before stabilizing at ~430mA, where it has remained.  The Filter Motor still appears to be stalled.  Options for recovery include toggling the Filter Motor on/off every orbit as we have done with success on NOAA-15/18 but are otherwise limited.  This effort is planned to start early next week, week of June 1. 


Topic:  Metop-B HIRS Filter Wheel Current Increase Event

Date/Time Issued: May 29, 2020 1836 UTC

Product(s) or Data Impacted:  HIRS Products

Date/Time of Initial Impact:  May 27, 2020 2240 UTC

Impacts on Users and Significance:  HIRS products are affected.

Date/Time of Expected End:  Unknown

Length of Outage:  Ongoing

Details/Specifics of Change:

The Metop-B HIRS Filter Wheel Motor current began showing signs of instability at approximately 2137 UTC on May 27, 2020.  At 2143 UTC it saw a large spike in current up to 476mA, then dropped to about 270mA, where it remained until 2240 UTC.  At that point, it spiked again to 418mA before ramping down to ~380mA where it has remained since.  The Filter Wheel Motor Temperature has since stabilized at ~30°.

This behavior is consistent with a filter wheel stall, but requires further investigation.  Options for recovery include putting the instrument into High Power mode but are otherwise limited.

Contact Information for Further Information:  ESPC Operations at ESPCOperations@noaa.gov at 301-817-3880 and/or other contacts as necessary

Web Site(s) for applicable information:  N/A

This message was sent to ESPC.Notification@noaa.gov.  You have been sent this and other notifications because you have opted in to receive it. If for any reason, you wish to unsubscribe, please contact ESPC Help Desk at ESPCOperations@noaa.gov or (301) 817-3880.  Please note: it make take up to two business days to process your unsubscribe request